How many kilograms of nickel silver containing 15% zinc and how many kilograms of nickel silver containing 30% zinc must be melted together to produce 704 kg of a new nickel silver containing 20% zinc?
Treat the percents given in this exercise as exact numbers, and work to three significant digits.

Answer by Alan3354(69443)

How many kilograms of nickel silver containing 15% zinc and how many kilograms of nickel silver containing 30% zinc must be melted together to produce 704 kg of a new nickel silver containing 20% zinc?
================================================
Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6